OpenStudy (anonymous):

solve 5x + 12 = x -4

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

5x+12=x-4 5x-x=-12-4 4x=-16 4x/4=-16/4 x=-4

OpenStudy (anonymous):

thanks so much you are the best there will be another one is a second